OpenStudy (anonymous):

Do you simplify trig ratios or no? For example: sin theta = -7,500/10,000

OpenStudy (anonymous):

Would I simplify that or keep it how it is?

OpenStudy (anonymous):

you can simplify

OpenStudy (anonymous):

its better to simplify otherwise your eyes will get hurt

Nnesha (nnesha):

\[\huge\rm \theta = \sin^{-1}\frac{ -7500 }{ 10000 }\] you need to find invs sin and yes you can simplify

OpenStudy (anonymous):

how do you make that big font?

Nnesha (nnesha):

``` \(\huge\color{green}{chocolates}\) ```

OpenStudy (anonymous):

ok thanks!

Nnesha (nnesha):

np o^_^o

Nnesha (nnesha):

BEST latex tutorial http://openstudy.com/study#/updates/52e12f56e4b0942cc9de719e

OpenStudy (anonymous):

i know how to latex...didnt know it works here lol

OpenStudy (anonymous):

usually they run on html

Nnesha (nnesha):

oh okay.

OpenStudy (anonymous):

thanks anyway!